OpenStudy (anonymous):

consider y''+ (1+ p(x) ) y= 0, where p is continuous in [x0, infinity ), lim p(x)= 0, integral of |p'(t)| dt < infinity. show that all solution of this DE are bounded in [x0, infinity ).

OpenStudy (anonymous):

i'm not sure

OpenStudy (freckles):

I don't think it helped

OpenStudy (anonymous):

can you please rewrite it i wanna make sure

OpenStudy (freckles):

I just solved for y... I don't know how that would help with p

OpenStudy (anonymous):

yeah i need to solve it for y

OpenStudy (freckles):

you will have to solve y in terms of the integral of p

OpenStudy (anonymous):

i didn't understand how did you get (y' v)' =0

OpenStudy (freckles):

\[y''v+v'y'=(vy')'\]

OpenStudy (freckles):

yep mistake I imagined y as y' for some reason

OpenStudy (anonymous):

yeah

OpenStudy (freckles):

so we want to show... \[x_0 \le y < \infty \\ \text{ given } \\ p \text{ is continuous on } [x_0,\infty) \\ \lim_{x \rightarrow \infty}p(x)=0 \\ \int\limits |p'(t)| dt <\infty \] doesn't the last condition mean p(t) is a constant function

OpenStudy (freckles):

and that would mean it would have to be p(t)=0 since as x goes to infinity p goes to 0

OpenStudy (freckles):

\[p(x)=0 \\ \text{ so the equation really is } y''=0\]

OpenStudy (anonymous):

y needs to be bonded by p

OpenStudy (freckles):

did I write the given parts above correctly?

OpenStudy (anonymous):

yes

OpenStudy (freckles):

if you didn't understand my reasoning for p(x) being 0 above \[\int\limits |p'(t)| dt =g(t) <\infty \\ \text{ means } g(t) \text{ is a constant number so } g(t)=C \in \mathbb{R}\\ \int\limits |p'(t)| dt=C \\ |p'(t)|=0 \\ p'(t)=0 \implies p(t)=K \in \mathbb{R} \\ \text{ and we are also given } \\ \lim_{x \rightarrow \infty} p(x)=0 \\ \text{ so we have } \\ \lim_{x \rightarrow \infty} K=0 \implies K=0 \\ \text{ so } p(x)=0\]

OpenStudy (anonymous):

i got it

OpenStudy (freckles):

maybe I'm missing something...

OpenStudy (anonymous):

so y< 0?

OpenStudy (freckles):

I don't know how you got that... I'm gonna mention some people that I think might be better at this than me: @jim_thompson5910 @sithsandgiggles @dan815 @ganeshie8 @IrishBoy123 @Kainui someone correct me on sith's name pretty please

OpenStudy (anonymous):

thanks a lot for your help